Let F =< 5x + 2, 2y + z, 10x + 10y > be a vector field in R3. Evaluate the following surface integral directly: XFX)DA Il vaš please answer will give rating thank you :)
Answer #1

Th F : <5x+ 2, 2y + 2,10% +107) ģCU,) = 0, VI 10-50-2v). Fu (UN) = <1,0,-5) řv (0,w) =<oil,-2). Tuxir =<s, 2, 1). = < SU +10-2 2] S-SU12 zo dudu: S-S012 du T 70 = J [ 70~7 블 ] 2 356 17 -s lol Jos도 2. 2 2 1 (I 3SO